OpenStudy (anonymous):

Find the slope of a line parallel to a line that contains the points (9, -3) and (-3, 8).

OpenStudy (anonymous):

when two line are parallel , their slopes are equal in order to determine the slope of the line containing points (x1,y1) and (x2,y2) use this formula (y2-y1)/(x2-x1)

OpenStudy (anonymous):

So, it would be (8-3)/(-3)/(9) ? @niksva

OpenStudy (anonymous):

m = {8- (-)3}/ (-3-9) = (8+3)/(-3-9)= 11/(-)12 = - 11/12